### v1

# Multiple stellar populations in MOCCA globular cluster models: Transient spatial over-concentration of pristine red giant stars driven by strong dynamical encounters

11institutetext: Nicolaus Copernicus Astronomical Center, Polish Academy of Sciences, ul. Bartycka 18, PL-00-716 Warsaw, Poland
11email: mig@camk.edu.pl; askar@camk.edu.pl
22institutetext: Faculty of Mathematics and Computer Science, A. Mickiewicz University, Uniwersytetu Poznańskiego 4, 61-614 Poznań, Poland 33institutetext: Korea Astronomy and Space Science Institute, Daejeon 34055, Republic of Korea
Multiple stellar populations in MOCCA globular cluster models: Transient spatial over-concentration of pristine red giant stars driven by strong dynamical encounters
M. Giersz1    A. Askar1    A Hypki1,2    J. Hong 3    G. Wiktorowicz1    L. Hellström1
(Accepted XXX. Received YYY; in original form ZZZ)
Recently, Leitinger et al. (2023) and Cadelano et al. (2024) have shown that, in some Milky Way globular clusters (GCs), pristine red giant branch (RGB) stars are more centrally concentrated than enriched RGB stars. This contradicts most multiple-population formation scenarios, which predict that the enriched population ($\rm 2P$) should initially be more concentrated than the pristine population ($\rm 1P$). We analyze a MOCCA GC model that exhibits a higher spatial concentration of $\rm 1P$ RGB stars than $\rm 2P$ RGB stars at 13 Gyr. The MOCCA models assume the asymptotic giant branch (AGB) pollution scenario, where $\rm 2P$ stars are initially more concentrated than $\rm 1P$ stars. Our results indicate that the observed spatial distributions of multiple populations, and possibly their kinematics, are significantly shaped by dynamical interactions. These interactions preferentially eject $\rm 2P$ RGB progenitors from the central regions, leading to a transient over-concentration of $\rm 1P$ RGB stars at late times. This effect is particularly relevant for GCs with present-day masses of a few $10^{5}\ \rm M_{\odot}$, which have retained only about 10–20% of their initial mass. Such clusters may appear dynamically young due to heating from a black hole subsystem, even if they have undergone significant mass loss and dynamical evolution. Additionally, the relatively small number of RGB stars in these clusters suggests that interpreting the spatial distributions of multiple populations solely from RGB stars may lead to biased conclusions about the overall distribution of $\rm 1P$ and $\rm 2P$.
The apparent over-concentration of the $\rm 1P$ relative to the $\rm 2P$ is likely a transient effect driven by the preferential removal of $\rm 2P$ RGB progenitors via strong dynamical encounters. MOCCA models of multiple stellar populations based on the AGB scenario may explain anomalous features observed in some Galactic GCs, such as NGC 3201 and NGC 6101.
Key Words.:
stellar dynamics – methods: numerical – globular clusters: evolution – stars: multiple stellar populations
1 Introduction
Globular clusters (GCs) were historically thought to host simple stellar populations with uniform chemical compositions. However, photometric and spectroscopic studies over the past few decades have revealed the presence of multiple stellar populations ($\rm MPOP$), primarily characterized by internal variations in light-element abundances (see Bastian & Lardo 2018; Gratton et al. 2019, and references therein). These populations include a pristine first population ($\rm 1P$), whose chemical composition resembles that of field stars with similar metallicities, and one or more enriched populations ($\rm 2P$), which exhibit signatures of high-temperature hydrogen burning (Milone & Marino 2022). The fraction of enriched $\rm 2P$ stars in Milky Way (MW) GCs strongly correlates with present-day cluster mass, ranging from approximately 40% in low-mass clusters to nearly 90% in the most massive ones (see Milone & Marino 2022, and references therein).
The formation of $\rm MPOP$ in GCs remains a topic of active debate, with several theoretical scenarios proposed to explain its origin. However, none of these scenarios fully account for all the observational features reported in the literature. Giersz et al. (2024) provides a concise overview of the most widely accepted formation mechanisms, particularly from the perspective of numerical simulations. The study also suggests that the post-formation migration of GCs may play a crucial role in shaping the properties of $\rm MPOP$. Most formation scenarios, including those that attribute $\rm MPOP$ to pollution from asymptotic giant branch (AGB) stars, predict that $\rm 2P$ should be more centrally concentrated than $\rm 1P$. Observations generally support this prediction, but exceptions exist such as the Galactic GCs NGC 3201 and NGC 6101, where $\rm 1P$ red giant branch (RGB) stars are found to be more centrally concentrated than $\rm 2P$ RGB stars (Leitinger et al. 2023). However, in the case of NGC 3201, Mehta et al. (2024), using RGB stars along with other evolved stars from the Horizontal Branch (HB), found no evidence of such over-concentration, as $\rm 1P$ and $\rm 2P$ appear to be well mixed.
In this short paper, we present results from a GC model simulated with the mocca code, where the $\rm 2P$ population initially has a much higher central concentration than the $\rm 1P$. However, by 13 Gyr, the $\rm 1P$ RGB stars become more spatially over-concentrated than the $\rm 2P$ RGB stars. The over-concentration of $\rm 1P$ RGB stars relative to $\rm 2P$ RGB stars may provide valuable insights into their formation scenarios and long-term evolution. It could also serve as an important constraint on the initial conditions of GC formation. Therefore, understanding the mechanisms driving this phenomenon and identifying the clusters where it may occur are crucial. In this study, we focus on the transient nature of $\rm 1P$ RGB over-concentration, offering a general explanation for its formation and the types of GCs in which it may be observed. A more detailed investigation will be presented in a forthcoming paper.
The paper is structured as follows: Section 2 provides a brief overview of the mocca code, Section 3 presents the simulation results for models exhibiting $\rm 1P$ over-concentration, and Section 4 interprets these findings and discusses their implications.
2 Method
This work is based on the numerical simulations performed with the mocca Monte Carlo code (Giersz 1998; Hypki & Giersz 2013; Giersz et al. 2013; Hypki et al. 2022, 2024; Giersz et al. 2024) mocca is an advanced code that performs full stellar and dynamical evolution of real size star clusters up to Hubble time. The implementation of stellar and binary evolution within the mocca code is based on the rapid population synthesis code bse code (Hurley et al. 2000, 2002) that has been strongly updated by Belloni et al. (2017, 2018), Banerjee et al. (2020), and Kamlah et al. (2022). Strong dynamical interactions in mocca are performed with fewbody code (Fregeau et al. 2004; Fregeau & Rasio 2007). mocca can follow the full dynamical and stellar evolution of $\rm MPOP$, allowing for delayed $\rm 2P$ formation relative to $\rm 1P$. In this scenario, $\rm 2P$ forms through the re-accretion of gas surrounding the newly formed cluster, which mixes with enriched material from the ejected envelopes of AGB stars. Additionally, mocca incorporates the effects of cluster migration, driven by the rapidly evolving gravitational potential of the early Galaxy and dynamical friction.
The stellar/binary evolution parameters used in the presented simulations are referred to as Model C in Kamlah et al. (2022). In short, the metallicity of both populations in all the simulated models was set to Z = 0.001. The updated treatment for the evolution of massive stars was used according to Tanikawa et al. (2020) together with improved treatment for mass loss due to stellar winds and the inclusion of pair and pulsational pair-instability supernova (Belczynski et al. 2016). The masses of black holes (BH) and neutron stars (NS) were determined according to the rapid supernovae prescriptions from Fryer et al. (2012). NS natal kicks were sampled from a Maxwelllian distribution with $\sigma=265$ km/s (Hobbs et al. 2005). However, for BHs, these natal kicks were reduced according to the mass fallback prescription (Belczynski et al. 2002; Fryer et al. 2012). The formation of neutron stars with negligible natal kicks through electron-capture or accretion-induced supernova was also enabled. Another feature of these models is the inclusion of gravitational wave (GW) recoil kicks whenever two BHs merge (Baker et al. 2008; Morawski et al. 2018). Low birth spins were assumed for BHs, with values uniformly sampled between 0 and 0.1 (Fuller & Ma 2019). The orientation of the BH spin relative to the binary orbit was randomly distributed (Morawski et al. 2018). The full list of changes in the recent mocca code together with their detailed description can be found in Giersz et al. (2024).
3 Results
Figure 1: Cumulative number distributions of RGB stars for $\rm 1P$ (black line) and $\rm 2P$ (red line) as a function of distance scaled by $R_{hl}$ for selected projected (2D) snapshots from 12 Gyr to 13.5 Gyr. $\rm 1P$ is more centrally concentrated than $\rm 2P$ only in the snapshot at 13 Gyr. The number distributions of both populations change significantly over time, and the over-concentration of $\rm 1P$ is a transient feature. The $R_{hl}$ values, snapshot times, and the number of RGB stars are provided in the insets of each panel.
Inspired by the work of Leitinger et al. (2023), who conducted a homogeneous analysis of $\rm MPOP$ in 28 Galactic GCs using a combination of HST photometry and wide-field, ground-based photometry of RGB stars, we examine a puzzling trend they identified. Their study found that in a few Galactic GCs, such as NGC 3201 and NGC 6101, $\rm 1P$ RGB stars are more concentrated within a few times the half-light radius than $\rm 2P$ RGB stars. This unexpected result is difficult to reconcile with current $\rm MPOP$ formation scenarios. Recently, Mehta et al. (2024) cast doubt on these findings, showing that in NGC 3201, when HB and AGB stars are included alongside RGB stars, both populations exhibit similar spatial distributions. However, Cadelano et al. (2024) confirmed the results of Leitinger et al. (2023) for NGC 3201, though they found that $\rm 2P$ RGB stars follow a bimodal distribution and are more concentrated within the half-light radius.
With these results in mind, we took a closer look at the $\rm MPOP$ simulations performed with the mocca code, as discussed in Giersz et al. (2024). We identified several models where $\rm 1P$ stars are more concentrated than $\rm 2P$ stars. From the work of Chen & Gnedin (2024), it is known that both NGC 6101 and NGC 3201 were formed ex-situ. Based on this, we selected a GC model with a capture time by the MW of 5.5 Gyr and an orbital migration time of 1 Gyr. In addition, we tried to select a model with observational properties similar to those of NGC 3201, specifically: cluster mass = 1.4464×$\times 10^{5}$ (1.93×$\times 10^{4}$) $M_{\odot}$, core radius $R_{c}$ = 1.52 (1.66) pc, observational core radius $R_{c,obs}$ = 3.32 (1.85) pc, half-light radius $R_{hl}$ = 5.98 (5.17) pc, half-mass radius $R_{h}$ = 8.71 (8.68) pc, tidal radius $R_{t}$ = 81.72 (82.34) pc, and the ratio between the $\rm 2P$ and $\rm 1P$ populations = 0.67 (0.46). The values for the mocca model were taken at 13 Gyr, while the values in parentheses represent the observed properties of NGC 3201 from Baumgardt et al. (2019) and Milone et al. (2017). The observational properties of the NGC 3201 are close to those of the model, but this does not mean that the model represents the actual evolution of the GC. It only serves as a possible example of GC evolution and we should not expect all cluster properties to be the same, in particular the $\rm 1P$ concentration range.The model is used to illustrate the possible causes of the observed $\rm 1P$ and $\rm 2P$ properties in NGC 3201.
Figure 1 shows the cumulative number distributions of RGB stars for $\rm 1P$ and $\rm 2P$ as a function of distance scaled by $R_{hl}$ at several selected snapshot times for a chosen mocca model. This model, whose 13 Gyr parameters were provided above, has the following initial conditions: number of objects $N_{\rm 1P}$ = 800000, $N_{\rm 2P}$ = 300000, and a Kroupa (2001) initial mass function (IMF), with $\rm 1P$ stars sampled between $\rm 0.08\ M_{\odot}$ and $\rm 150\ M_{\odot}$, and $\rm 2P$ stars between $\rm 0.08\ M_{\odot}$ and $\rm 20\ M_{\odot}$. The metallicities are $\rm Z_{1P}$ = 0.001 and $\rm Z_{2P}$ = 0.001, with binary fractions $fb_{\rm 1P}$ = 0.95 and $fb_{\rm 2P}$ = 0.95111Initial binary parameters were distributed according to Belloni et al. (2017). The Galactocentric distance is $R_{g}$ = 1.0 kpc, and the ratio of half-mass radii is $R_{h_{2P}}/R_{h_{1P}}$ = 0.1. $\rm 1P$ and $\rm 2P$ are separately in virial equilibrium, with $\rm 1P$ being tidally filling (with $R_{h_{1P}}=9.03$ pc and $R_{t}/R_{h_{1P}}=3.66$). The initial King (1966) model concentration parameter is $\rm W_{0}=3$ for $\rm 1P$ and $7$ for $\rm 2P$. Over 1 Gyr following 5.5 Gyr of evolution, the cluster migrates along its circular orbit from 1 kpc to 5 kpc, mimicking the capture of a GC during a dwarf galaxy merger with the MW.
[…]
Figure 1 clearly shows that $\rm 1P$ is more centrally concentrated than $\rm 2P$ only in the snapshot at 13 Gyr. In other snapshots, both populations are well mixed, or $\rm 2P$ is more concentrated than $\rm 1P$. The number distributions of both populations vary significantly over time, indicating that the over-concentration of $\rm 1P$ is a transient feature. The magnitude of $\rm 1P$ over-concentration at 13 Gyr is not as pronounced as that observed for NGC 3201 by Leitinger et al. (2023) and Cadelano et al. (2024). This is not unexpected, as the presented model is only a rough approximation of an actual GC. As discussed in Section 4, the extent of $\rm 1P$ over-concentration depends on the evolutionary details of RGB progenitors and the initial conditions of the cluster model.
[…]
In the next section (Section 4), we further analyze the properties of models that explain why $\rm 1P$ is more concentrated than $\rm 2P$, why the over-concentration of $\rm 1P$ is observed only in RGB stars, and why this feature appears to be transient.
4 Discussion and Conclusions
4.1 Black hole subsystem and cluster migration
By analyzing mocca simulations with the $\rm MPOP$ implementation, as described in Giersz et al. (2024), and focusing on models where $\rm 1P$ is more centrally concentrated than $\rm 2P$, we draw the following conclusions. The over-concentration of $\rm 1P$ is observed in GC models that have lost 80–90% of their initial mass, leading to present-day (11–13 Gyr) masses of approximately $\rm 10^{5}\ M_{\odot}$. Additionally, these clusters retain a BHS comprising up to a few hundred stellar-mass BHs at present-day. In clusters where the present-day mass is only 10–20% of the initial mass and the BHS accounts for approximately 5% of the total cluster mass, the mass distribution of $\rm 1P$ becomes more centrally concentrated than that of $\rm 2P$. This is because the majority of BHs originate from the evolution of massive $\rm 1P$ stars, as the IMF for $\rm 1P$ extends up to $\rm 150\ M_{\odot}$, whereas for $\rm 2P$ stars, the IMF is limited to a maximum mass of $\rm 20\ M_{\odot}$, significantly reducing BH formation from $\rm 2P$ stars.
The GC model described in this paper initially retains nearly 1000 stellar-mass BHs at around 50 Myr, originating from the evolution of massive $\rm 1P$ progenitors. At this time, the 10% and 50% Lagrangian radii for these BHs are approximately 4.8 pc and 10.8 pc, respectively. As the most massive objects in the cluster, BHs begin to segregate toward the center. The delayed inclusion of the centrally concentrated $\rm 2P$ population between 50 and 100 Myr accelerates this segregation. By 500 Myr, the 10% and 50% Lagrangian radii for BHs contract to about 1 pc and 5.4 pc, respectively, continuing to shrink until around 2 Gyr. The BHS, consisting of both single and binary BHs, strongly influences the cluster’s dynamical evolution through interactions with other stars (Mackey et al. 2008; Breen & Heggie 2013; Sippel & Hurley 2013; Morscher et al. 2015; Arca Sedda et al. 2018; Kremer et al. 2019). Since the $\rm 2P$ population is initially more centrally concentrated in this model, $\rm 2P$ stars are more likely to undergo strong encounters with $\rm 1P$ BHs and other $\rm 2P$ stars. These interactions can scatter $\rm 2P$ stars onto wider orbits or eject them from the cluster entirely (see Section 4.2).
For both NGC 3201 and NGC 6101, where $\rm 1P$ RGB stars appear more centrally concentrated than $\rm 2P$ RGB stars, several studies suggest that these GCs likely harbor a significant population of stellar-mass BHs (Peuten et al. 2016; Askar et al. 2018; Kremer et al. 2018; Vitral et al. 2022). Furthermore, three stellar-mass BHs have been identified in NGC 3201 through radial velocity variations using MUSE spectroscopic observations (Giesers et al. 2018, 2019). Therefore, it is possible that the possible presence of a BHS has played a role in shaping the spatial distribution of $\rm 1P$ and $\rm 2P$ RGB stars in these GC.
The GC NGC 3201 is likely an ex-situ cluster that was accreted by the MW along with a dwarf galaxy (Chen & Gnedin 2024). It is characterized by large $R_{h}$ and $R_{t}$ radii. During the early stages of its evolution, it likely experienced a strong tidal field within its parent galaxy, and it can be assumed that it was initially tidally filling, with small $R_{h}$ and $R_{t}$ radii. After being captured by the MW, the tidal field weakened considerably, causing the cluster to become tidally underfilling, allowing it to expand freely within its $R_{t}$ and significantly increase its half-mass relaxation time. For these reasons, we selected a cluster model that initially forms near the galactic center with an initial mass of approximately $10^{6}M_{\odot}$. Over a few Gyr, the cluster migrates to a larger Galactocentric distance (see Section 3 for details). This model not only reproduces similar observational properties to NGC 3201 but also yields a comparable ratio of $\rm 2P$ to $\rm 1P$ stars. Due to the cluster’s relatively low present-day mass, the number of RGB stars is relatively small, on the order of several hundred. As we show below, such a small number of RGB stars plays a crucial role in the over-concentration of $\rm 1P$ relative to $\rm 2P$ RGB stars.
Figure 3: The lower panel shows the escape time versus radial position for dynamically ejected present-day RGB progenitor stars with initial masses between 0.81 and 0.83 M. Each point represents a present-day RGB progenitor that escaped following a strong three-body (binary–single) or four-body (binary–binary) encounter. The x-axis indicates the escape time, while the y-axis shows the radial position within the cluster at which the interaction that led to the escape occurred. Superimposed on the scatter plot are colored lines tracing the evolution of the cluster’s core radius ($R_{c}$, grey), half-mass radius ($R_{h}$, dark blue), half-light radius ($R_{hl}$, light blue) and tidal radius ($R_{t}$, brown). Violet symbols denote $\rm 1P$ escapers and green symbols denote second-population ($\rm 2P$) escapers; filled symbols indicate single escapers, whereas open symbols correspond to binary escapers. In total, there are 150 $\rm 1P$ escapers (of which 34 are in binaries) and 402 $\rm 2P$ escapers (with 76 in binaries). The red dashed vertical line marks the 13 Gyr snapshot, at which the over-concentration of $\rm 1P$ RGB progenitors is evident. The top panel displays a histogram of the escape times, binned in intervals of 200 Myr for both populations, highlighting that a greater number of $\rm 2P$ RGB progenitors were removed from the cluster due to strong encounters.
4.2 Preferential ejection of enriched RGB progenitors by strong dynamical encounters
To explain the spatial over-concentration of $\rm 1P$ RGB stars at 13 Gyr, we first identified stars that evolve into RGBs at this epoch. These stars are near the turn-off mass, which is 0.8154 $\rm M_{\odot}$ for $\rm Z=0.001$ and had zero-age MS mass of approximately 0.83 $\rm M_{\odot}$. Assuming that all stars in the initial model (or at $\sim$100 Myr, when $\rm 2P$ is fully added) with masses between 0.81 and 0.83 $\rm M_{\odot}$ will have evolved into RGB stars by 13 Gyr, we define this mass range as the present-day RGB progenitors and examined those that escaped before 13 Gyr.
For $\rm 1P$ stars, we counted all escaping RGB progenitors within the 0.81–0.83 $\rm M_{\odot}$ mass range, regardless of the cause of escape, and found that 5048 $\rm 1P$ RGB progenitors had escaped the cluster. Similarly, for $\rm 2P$ stars, 1167 RGB progenitors had escaped. This significant difference is expected, as $\rm 1P$ was initially tidally filling. However, when considering only escapers due to strong dynamical interactions (binary-single or binary-binary encounters), the numbers shift dramatically: $\rm 1P$ RGB progenitor escapers drop to 150, while $\rm 2P$ increases to 403. Thus, $\rm 2P$ RGB progenitors are 2.7 times more likely to be dynamically ejected than $\rm 1P$ progenitors, consistent with their initial central concentration.
We find that $\rm 2P$ RGB progenitors are more frequently ejected from the inner cluster regions due to strong dynamical encounters. Matching escaping stars with interaction data shows that most present-day $\rm 2P$ RGB progenitors were ejected between 2 and 6 Gyr, primarily from within a few times the half-light radius ($R_{hl}$). After 6 Gyr, following the cluster’s migration to a larger Galactocentric radius, the ejection rate decreases, though $\rm 2P$ RGB progenitors continue to be preferentially removed. This trend is illustrated in Fig. 3, where the lower panel shows that most $\rm 2P$ escapers originate from the inner regions, especially before 6 Gyr, while the top panel histogram highlights the higher escape rate of $\rm 2P$ RGB progenitors compared to $\rm 1P$. These results confirm that strong dynamical encounters selectively remove $\rm 2P$ RGB progenitors, contributing to the observed over-concentration of $\rm 1P$ RGB stars at 13 Gyr.
To quantify this effect, we define the ratio:
$R=\frac{N_{\text{RGB,in\_cluster}}}{N_{\text{RGB,progenitor\_dyn\_escapers}}},$
where $N_{\text{RGB,in\_cluster}}$ is the number of RGB stars present in the cluster at a given snapshot, and $N_{\text{RGB,progenitor\_dyn\_escapers}}$ is the cumulative number of RGB progenitors lost to strong encounters by that time. For $\rm 2P$ at 13 Gyr, we find $R\approx 1.06$ (430/402), meaning that the number of remaining RGB stars is nearly equal to those lost through strong encounters. At earlier times, such as 12.5 Gyr, this ratio is higher ($R\approx 1.13$), as fewer progenitors have escaped relative to in-cluster RGB stars. By 13.5 Gyr, the ratio returns to $R\approx 1.13$, indicating that the over-concentration of $\rm 1P$ relative to $\rm 2P$ RGB stars is most pronounced at 13 Gyr, when $R$ is closest to unity. This is further amplified by the small-number statistics of RGB stars relative to the total stellar population. At 13 Gyr, the number of $\rm 2P$ RGB stars in the cluster (430) is nearly equal to the number of progenitors ejected due to strong encounters (402). This near-parity magnifies the impact of missing $\rm 2P$ RGB progenitors. At earlier times (12.5 Gyr), a larger fraction of $\rm 2P$ RGB stars remain in the cluster ($R>1.1$), mitigating this effect. However, at 13 Gyr, when $R\approx 1.06$, the influence of escaped progenitors dominates, leading to the over-concentration of $\rm 1P$ RGB stars. By 13.5 Gyr, the ratio increases again, erasing this over-concentration.
This analysis highlights the distinct dynamical evolution of $\rm 2P$ RGB progenitors, which are preferentially ejected through strong interactions, particularly from the cluster’s inner regions. The interplay of small-number statistics and dynamical encounters thus plays a crucial role in shaping the observed RGB spatial distribution.
4.3 Conclusion
Our study suggests that the spatial distribution and, potentially, the kinematic properties of $\rm MPOP$ may depend on the type of stars observed. This effect is particularly relevant for GCs with present-day masses of a few $10^{5}\ \rm M_{\odot}$, which have retained only 10–20% of their initial mass. Despite the significant mass loss, such clusters may appear dynamically young due to BHS heating and migration to larger Galactocentric distances. When observations are limited to RGB stars, small-number statistics and dynamical interactions can distort their spatial distribution, leading to biased conclusions about the overall distribution of $\rm MPOP$. The over-concentration of $\rm 1P$ RGB stars relative to $\rm 2P$ RGB stars appears to be a transient effect.
Confirming the findings of Leitinger et al. (2023) and Cadelano et al. (2024), that in some MW GCs, $\rm 1P$ RGB stars are more centrally concentrated than $\rm 2P$ RGB stars, requires observations of MS stars. Our mocca simulations show no $\rm 1P$ over-concentration relative to $\rm 2P$ among MS stars. Verifying this observationally would be a crucial step toward understanding $\rm MPOP$ properties and supporting the AGB scenario of $\rm MPOP$ formation. Additionally, our results align with the kinematic analysis by Cadelano et al. (2024), which suggests that $\rm 2P$ stars initially formed more centrally concentrated than $\rm 1P$ stars, influencing their present-day spatial and kinematic distributions.
We have shown that the observed over-concentration of $\rm 1P$ relative to $\rm 2P$ can be explained within the AGB scenario, provided certain internal and external conditions are met. This effect appears to be transient, driven by fluctuations in the number and spatial distribution of RGB stars. Further analysis is needed to confirm these results, particularly regarding binary fraction distributions and velocity anisotropy in $\rm 1P$ and $\rm 2P$. In an upcoming study, we will conduct a more detailed comparison of mocca model results with observations from Leitinger et al. (2024) and Cadelano et al. (2024) to further investigate these trends.
Acknowledgments
MG, AH, GW, LH were supported by the Polish National Science Center (NCN) through the grant 2021/41/B/ST9/01191. AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Data Availability
The data for the mocca model described in this paper, including the 13 Gyr snapshot, is available at https://doi.org/10.5281/zenodo.14908094. Additional data can be provided upon request to the corresponding author.